A plane wave front is incident on a surface of glass slab of R.I. 1.5 making an angle 60°.with it's surface. Find angle of refraction.

Grade:12

1 Answers

rajasekhar reddy
5 years ago
U(mu)=sin i/sin r
1.5=sin60/sin r
3/2=(root3/2)/sinr
sin r=((root3/2)*2)/3
sin r =root3/(root3*root3)
sin r=1/root3
r=sin inverse1/root3
r=0.6154